### v2

# Pitfalls and Outlooks in Using COMET

Abstract
Since its introduction, the COMET metric has blazed a trail in the machine translation community, given its strong correlation with human judgements of translation quality. Its success stems from being a modified pre-trained multilingual model finetuned for quality assessment. However, it being a machine learning model also gives rise to a new set of pitfalls that may not be widely known. We investigate these unexpected behaviours from three aspects: 1) technical: obsolete software versions and compute precision; 2) data: empty content, language mismatch, and translationese at test time as well as distribution and domain biases in training; 3) usage and reporting: multi-reference support and model referencing in the literature. All of these problems imply that COMET scores is not comparable between papers or even technical setups and we put forward our perspective on fixing each issue. Furthermore, we release the SacreCOMET package that can generate a signature for the software and model configuration as well as an appropriate citation. The goal of this work is to help the community make more sound use of the COMET metric.
mdframedlisting backgroundcolor=black!10, linecolor=white, innertopmargin=-2mm, innerbottommargin=0mm, innerrightmargin=0mm, innerleftmargin=0mm,
Pitfalls and Outlooks in Using COMET
Vilém Zouhar 1 Pinzhen Chen 2 Tsz Kin Lam2 Nikita Moghe2 Barry Haddow2 1ETH Zurich 2University of Edinburgh vzouhar@ethz.ch {pinzhen.chen,tlam,nikita.moghe,bhaddow}@ed.ac.uk
1 Introduction
Automated metrics provide a cheap and scalable way of evaluating and benchmarking NLP models. In machine translation (MT), the evaluation protocol has moved from string matching metrics (BLEU, TER, chrF, inter alia; Papineni et al., 2002; Snover et al., 2006; Popović, 2015) to trained metrics based on neural networks (Shimanaka et al., 2018; Takahashi et al., 2020; Rei et al., 2020a; Sellam et al., 2020) with COMET being the most widely used option in the community. The trained metrics have been shown to correlate much better with human judgement (Freitag et al., 2021, 2022b, 2023), making them more reliable in assessing or estimating translation quality and ranking translation systems.
Nonetheless, the solution to translation evaluation is yet to be perfected. One problem is the misuse and misunderstanding of the metric. Previously, Post (2018) showed that different usages and implementations of BLEU, e.g. tokenization and smoothing, can lead to vast inconsistencies in reported scores. We suspect that sensitivities can arise from several stages of using neural metrics too—e.g. obtaining COMET involves running a particular checkpoint with the COMET software on some particular hardware. This process might be sensitive to misconfigurations in each stage, resulting in unexpected or nonreproducible behaviours. Furthermore, trained MT metrics like COMET often fit on limited data intended to be valid translations, leading to reduced robustness owing to its nature as an optimized machine learning model.
In this work, we
list nine problems covering technical issues, data bias, and model reporting;
we show that inconsistent use of COMET leads to scores non-comparable across papers or setups;
release SacreCOMET package for better reporting and reproducibility.
2 Background and Setup
Metric background.
Human judgements for translation quality come from annotation campaigns, where translations are evaluated with some annotation protocol. Historically for WMT, the protocol was a variant of direct assessment (DA; Graham et al., 2013) which has annotators providing a number from 0 (lowest) to 100 (highest) as the overall segment quality. This has been subsequently replaced by MQM and ESA protocols (Lommel et al., 2014; Kocmi et al., 2024c), though DA remains the most abundant data source for neural metric training.
Automated metrics aim to yield scores that correlate with human judgements of translations. Most metric scores are computed at the segment level and then aggregated at the system level. The evaluation of metrics is done with respect to the human judgements.
COMET models.
Metrics such as chrF or BLEU are heuristic algorithms that match n-grams between the translation and the reference to compute a score. In contrast, COMET is a machine learning model fine-tuned from a pre-trained multilingual language model, e.g. XLM-R (Conneau et al., 2020), with an additional regression head. A reference-based COMET model learns to regress from a tuple of [source, hypothesis, reference] to the human judgement score (from previous evaluation campaigns) at the segment level. The quality estimation version of COMET is prepared by omitting the reference in the input.
Most issues in this work are demonstrated using two COMET checkpoints unless noted otherwise: the reference-based COMET and the reference-free COMET (Rei et al., 2022a).111huggingface.co/Unbabel/wmt22-comet-da,222huggingface.co/Unbabel/wmt22-cometkiwi-da Both metrics output to a normalized range between 0 and 1. The COMET framework unbabel-comet is of version 2.2.2 except when we test how different software versions affect COMET scores.
Data setup.
3 Problems
In this section, we identify and test nine possible pitfalls or curious behaviours with COMET which are not all well-studied. In three groups, these are:
Technicality: obsolete Python and COMET software versions as well as compute precisions could lead to inaccurate score computation.
Training and test data: COMET as a neural metric, might be derailed by empty hypotheses, language mismatch, and translationese at test time. It may also follow the training data biases.
Tool usage and score interpretation: it has no defined way of equipping multiple references when available which leaves room for research. From a bibliometric perspective, we reveal that some literature omits a clear reference to the checkpoint version or citation.
In addition, we discuss some final issues that need more attention from the community.
| Python | 3.7.16 | 3.8.11 | 3.12.4 |
| unbabel-comet | 1.1.2 | 2.2.2 | 2.2.2 |
| EnDe | 0.796 | 0.837 | 0.837 |
| EnZh | 0.911 | 0.862 | 0.862 |
| DeEn | 0.851 | 0.855 | 0.855 |
| ZhEn | 0.795 | 0.803 | 0.803 |
3.1 Software versions [technical]
The official installation of the COMET package requires Python 3.8 or above.333github.com/Unbabel/COMET 332dfb0 as of Aug 2024. We demonstrate that neglecting this would lead to unexpected scores because the same COMET checkpoint can produce vastly different scores with previous COMET framework versions that are no longer supported.
Under several Python versions, executing the following code leads to different COMET package (unbabel-comet) versions being installed. Running the framework for translation evaluation will subsequently result in false conclusions as shown in Table 1’s evaluation on WMT23 tests. The direct cause is that the discontinued Python 3.7 supports only up to unbabel-comet 1.1.2. Nonetheless, we caution that the underlying factor is the version of unbabel-comet rather than Python.
[style=mdframedlisting]
Recommendation.
Updating both Python and unbabel-comet to the latest version is helpful and reporting the version of the toolkit can aid reproducibility. The package version should also be reported.
| Precision | COMET | MAE | Acc | sec | ||
| EnDe | GPU FP32 | 0.822 | 10.4 | 0.274 | 0.885 | 113 |
| FP16 | 0.822 | 10.4 | 0.274 | 0.885 | 55 | |
| CPU FP32 | 0.822 | 10.4 | 0.274 | 0.885 | 2262 | |
| FP16 | 0.822 | 10.4 | 0.274 | 0.885 | 2403 | |
| QINT8 | 0.852 | 10.8 | 0.109 | 0.385 | 1856 | |
| DeEn | GPU FP32 | 0.841 | 9.98 | 0.296 | 0.901 | 87 |
| FP16 | 0.841 | 9.98 | 0.296 | 0.901 | 48 | |
| CPU FP32 | 0.841 | 9.98 | 0.296 | 0.901 | 1674 | |
| FP16 | 0.841 | 9.99 | 0.296 | 0.901 | 1758 | |
| QINT8 | 0.860 | 10.7 | 0.164 | 0.516 | 1249 | |
| EnZh | GPU FP32 | 0.842 | 11.7 | 0.290 | 0.933 | 111 |
| FP16 | 0.842 | 11.7 | 0.290 | 0.933 | 102 | |
| CPU FP32 | 0.842 | 11.7 | 0.290 | 0.933 | 1751 | |
| FP16 | 0.842 | 11.7 | 0.290 | 0.933 | 1710 | |
| QINT8 | 0.881 | 13.2 | 0.031 | 0.608 | 1258 | |
| ZhEn | GPU FP32 | 0.799 | 9.95 | 0.153 | 0.717 | 113 |
| FP16 | 0.799 | 9.95 | 0.153 | 0.725 | 86 | |
| CPU FP32 | 0.799 | 9.95 | 0.153 | 0.717 | 1936 | |
| FP16 | 0.799 | 9.95 | 0.153 | 0.717 | 1995 | |
| QINT8 | 0.872 | 10.5 | 0.081 | 0.475 | 1351 |
3.2 Numerical precision [technical]
Model quantization represents a model using lower-numerical precision data types so that the model consumes less memory and model passes can be computed faster. Such improvement in inference is directly beneficial to deployment efficiency; it is also useful in other complex procedures involving COMET scoring, such as data filtering, re-ranking, and Minimum Bayes Risk (MBR) decoding (Kumar and Byrne, 2004).
Despite the aforementioned advantages, model quantization is not a feature supported by the current COMET framework by default. We thus make minimal modifications to the software and investigate the effect of numerical precision on COMET scores on both CPU (FP32, FP16, and QINT8) and GPU (FP32 and FP16). For CPU inference with FP16 or dynamic QINT8, we apply the quantization module torch.ao.quantization from PyTorch before model.predict(). When using GPU-FP16, we first load the model weights to FP32. After that, we call .half() to explicitly convert the weights of the encoder, i.e., XLMRobertaModel() from transformers. This is because we discovered that loading the weights directly in FP16 would still result in FP32 precision.
| COMET | COMET | |||||||
| EnDe | EnZh | DeEn | ZhEn | EnDe | EnZh | DeEn | ZhEn | |
| Baseline (Online-A) | 0.837 | 0.862 | 0.855 | 0.803 | 0.800 | 0.791 | 0.794 | 0.787 |
| All empty | 0.335 | 0.392 | 0.353 | 0.374 | 0.315 | 0.319 | 0.537 | 0.467 |
| Random sent. | 0.373 | 0.434 | 0.334 | 0.350 | 0.333 | 0.341 | 0.447 | 0.391 |
| Random sent. w/ words shuffled | 0.244 | 0.419 | 0.264 | 0.347 | 0.232 | 0.325 | 0.307 | 0.385 |
We use NVIDIA GeForce RTX 3090 for GPU inference and a batch size of 8 in all settings (in practice a quantized model makes room for a larger batch size). Table 2 summarises the effect of numerical precision. In addition to reporting COMET scores, we also report 1) inference time in seconds (sec) as an efficiency measure; and 2) segment-level mean absolute error (MAE), segment-level Kendall’s tau-c (), and system-level pairwise accuracy (Acc). Everything is compared to the human DA scores either on segment- or system-level. Technically, Kendall’s calculates rank correlation on an ordinal scale with adjustments for ties and pairwise accuracy computes the proportion of system pairs that have the same ordering by a metric as by humans.
Our results show that there is no meaningful difference between FP32 and FP16 in both CPU and GPU devices up to 3 significant figures. On GPU, FP16 is about 30% faster in time, but unsurprisingly it does not provide any speed-up on CPU. Interestingly, on the CPU, dynamic QINT8 gives systematically higher COMET scores and shorter running time than FP32 and FP16. However, the much lower and pairwise accuracy indicate the lack of reliability at this precision. In addition to precision, we explored the effect of batch size and the choice of GPU or CPU during inference with results listed Appendix A. Whilst there are some fluctuations, they are mostly negligible. However, lower precision allows for higher batch size which usually directly corresponds to speed-up.
Recommendation.
If GPU is available, it is feasible to run COMET with FP16 with a larger batch size for much faster inference without any quality loss. Otherwise, FP32 should be used.
3.3 Empty hypothesis [data]
An empty translation (string of length 0) gets penalized heavily by string-based metrics because an empty string has zero surface overlap with the reference. However, neural metrics provide no such guarantee. We show that COMET assigns a positive instance-level score even if the hypothesis is an empty string as corroborated by Lo et al. (2023). Problematically, this score can even occasionally be higher than the score for a genuine system translation.
In Table 3, we list COMET scores for system Online-A’s hypotheses at WMT23 and a file full of empty lines. Furthermore, we compare them with completely incorrect translations to explicate the score magnitude in two ways. 1) Random sent.: we shuffle at the sentence level WMT22’s reference files in the respective translation directions. This provides us with high-quality human-written sentences. We sub-sample or over-sample if the number of lines in WMT22 is larger or smaller than the WMT23 size. 2) Random sent. w/ words shuffled: we further shuffle the words at each line in the sentence-shuffled files, generating nonsensical sentences. Sentence-shuffled hypotheses can be seen as fluent but extremely inadequate sentences whereas word-shuffled sentences are neither fluent nor adequate.
We observe that sentence-shuffled hypotheses attain comparable scores to empty ones, but word-shuffled hypotheses have the lowest scores across the majority of the translation directions. Empty and shuffled hypotheses, despite having much lower COMET than the valid translations, would not be assigned zero scores by COMET or COMET, showing that COMET is more lenient than string overlap-based metrics in penalizing such irregularities.
We then count the number of empty lines that score better than a translation from Online-A in Table 4. We observe roughly 0.25% such cases for most translation directions except for DeEn’s COMET score situating at 1.45%. Further, in Figure 1 we plot the distributions of COMET scores for ZhEn’s empty and genuine translations with other translation directions in Appendix B. Noticeable overlaps are observed for COMET when translating into English.
Adhering to the guidelines of the DA protocol, this is not the correct behaviour because an empty hypothesis should receive a score of 0, to match the category no meaning preserved. This, however, is not surprising with COMET, because it has likely not seen empty hypotheses during training that would have received a score of 0 from a human annotator. Finally, since the distributions of empty hypotheses are nowhere close to a single vertical bar, it implies that segment-level COMET scores oddly hinge on the source sentence, as noted by Sun et al. (2020).
| translation < empty | ||
| COMET | COMET | |
| EnDe | 0 / 558 | 2 / 558 |
| EnRu | 1 / 2075 | 0 / 2075 |
| EnUk | 2 / 2075 | 1 / 2075 |
| EnZh | 6 / 2075 | 2 / 2075 |
| DeEn | 1 / 550 | 8 / 550 |
| RuEn | 5 / 1724 | 6 / 1724 |
| UkEn | 1 / 1827 | 5 / 1827 |
| ZhEn | 5 / 1977 | 1 / 1977 |
Recommendation.
Force empty hypotheses to have 0 scores before aggregating. Also, a string-based metric like BLEU or chrF should be used to catch similarly malformed hypotheses.
3.4 Hypothesis language mismatch [data]
String overlap-based metrics can also score a hypothesis in a language different from the reference almost zero. However, even for the reference-based COMET, there is no explicit way to enforce the intended target language given it is an embedding-based metric. In the field of translation and translation evaluation, this poses an increasingly pronounced problem, especially for multilingual translation models as well as the recent large language models, in which the generated language cannot be as easily controlled as it could be in encoder-decoder models (Zhang et al., 2023).
We conduct experiments to understand if translation outputs in an incorrect language leave an impact on the score, and whether different mismatching languages can lead to distinct patterns. We use the translation directions EnRu, EnUk, and EnZh in WMT23 which share the same English source input. Having Online-A’s output in all three directions, we substitute hypotheses in a particular translation direction with those from another direction. A similar hypothesis was presented in Amrhein et al. (2022) which suggested that COMET metrics are not robust to hypothesis language mismatch. Our experiment setup offers a more detailed evaluation setup than their contrastive setup.
Table 5 presents COMET scores for translations in correct and incorrect languages, as well as empty lines and random sentences in the correct language as “baselines”, deemed as completely wrong translations. The pattern shows that when the hypothesis is in a language distant from the reference, COMET score declines much more than when the hypothesis is in a more similar language. A more concerning issue is that even when the hypotheses’ language is completely wrong, the resulting COMET score can still be vastly higher than empty hypotheses or random sentences in the correct language. We omit COMET because it does not have a mechanism to read a reference (language) making it inherently incapable of distinguishing output languages.
Recommendation.
Run language identification and set hypotheses in an unexpected language to have 0 COMET score before aggregating them system-level. Also, check with a string overlap-based metric like BLEU or chrF.
| Mismatch () Hyp. lang () | EnRu | EnUk | EnZh | |||
| Uk | Zh | Ru | Zh | Ru | Uk | |
| 100% | 0.797 | 0.536 | 0.807 | 0.540 | 0.655 | 0.644 |
| 0% | 0.853 | 0.832 | 0.862 | |||
| All empty | 0.316 | 0.329 | 0.391 | |||
| Random sent. | 0.463 | 0.472 | 0.435 | |||
3.5 Score distribution bias [data]
As the COMET metric is a machine learning model trained on human ratings of existing machine translations, it inherits many properties of statistical learning such as data (output score) distribution bias. The yearly WMT shared task receives submissions with varying quality and potentially varying quality ranges for different translation directions per year (from Koehn and Monz, 2006 to Freitag et al., 2023). This is attributed to diverse factors: the availability of data, source-target language similarity, the level of interest in languages, etc. The gap in translation quality will then propagate into skewed human judgement scores across translation directions. When a single COMET model learns to score all translation directions, it can overfit the score distribution w.r.t. a translation direction in addition to the quality of a translation hypothesis.
| Lang | Score |
| De En | 0.754 |
| Ps En | 0.670 |
| Is En | 0.724 |
| Pl En | 0.761 |
| Ru En | 0.771 |
| Ja En | 0.663 |
| Ta En | 0.655 |
| Zh En | 0.743 |
| Ha En | 0.641 |
| Km En | 0.659 |
| Lt En | 0.726 |
| Cs En | 0.740 |
| Gu En | 0.575 |
| Kk En | 0.649 |
| Iu En | 0.724 |
| Fi En | 0.719 |
| Lang | Score |
| En Is | 0.666 |
| En Lt | 0.600 |
| En Ru | 0.765 |
| En Iu | 0.720 |
| En Ha | 0.768 |
| En Ja | 0.745 |
| En Pl | 0.706 |
| En Gu | 0.514 |
| En Cs | 0.767 |
| En Zh | 0.775 |
| En Fi | 0.616 |
| En Ta | 0.709 |
| En De | 0.841 |
| En Kk | 0.574 |
| Lang | Score |
| Hi Bn | 0.910 |
| De Fr | 0.792 |
| Fr De | 0.834 |
| Zu Xh | 0.639 |
| De Cs | 0.510 |
| Xh Zu | 0.574 |
| Bn Hi | 0.770 |
We first verify this in Table 6 which shows that WMT translation directions are associated with vastly different human DA scores (from 0.51 to 0.91). Empirically, we illustrate this issue using two high-resource directions EnDe and EnZh. As shown in Figure 2, we keep either the top 75% or bottom 75% of all scored translations to alter the score distribution for each translation direction, simulating the scenario where low- and high-performing system submissions are received for different directions. We then train different COMET models on the data before and after alteration as per Figure 2. Finally, we evaluate those checkpoints on the same test set and report results in Table 7.
As expected, for both EnDe and EnZh, training on the top or bottom-scoring data would lead to increased or decreased COMET scores on the same set of hypotheses. In addition, we observe that altering scores in a particular translation direction would not affect scores in another direction. For example, removing the bottom 25% EnDe scores “improves” the test-time score from 0.770 to 0.790 whilst EnZh remains unaffected.
| Training data | EnDe | EnZh | |
| All | 0.770 | 0.770 | |
| Top-75% | of EnDe | 0.790 | 0.770 |
| Bot-75% | of EnDe | 0.765 | 0.764 |
| Top-75% | of EnZh | 0.783 | 0.789 |
| Bot-75% | of EnZh | 0.772 | 0.751 |
Implication.
A trivial conclusion is that COMET scores for different translation directions are not comparable. Nevertheless, we caution that the same phenomenon could happen for other features such as the domain, output style, etc. After all, although z-score re-scaling could mitigate this problem, we still encounter two issues. First, it has not been a common practice since WMT22 and it would further contribute to non-objective scores (Knowles, 2021); second, while z-scoring is straightforward for the translation direction, other latent, language-agnostic biases still exist.
3.6 Domain bias [data]
Zouhar et al. (2024) found that neural metrics like COMET are biased towards particular domains, manifested by worse test performance on unseen domains. Taking inspiration from their work and our discussions on “latent biases” in the previous section, we now raise a question—can we create adversarial hypotheses in test time to exploit the domain bias in training time? Specifically, different domains in the training data can be associated with different score ranges. Hence, by pretending that a test translation is in a particular domain, it might be possible to manipulate its COMET score.
In our experimental setting, to make it explicit to COMET during training, we prepend the target translation with a tag of its domain—in our case, the year the WMT data originated. Note that in each iteration of WMT, systems get higher overall DA scores (e.g. 0.721 for 2019 and 0.749 for 2021). Table 8 illustrates our setup: during training, we tag the scored translation data with its year; during testing, we trial various year prefixes to understand the effect.
| Train | 2020 Fire prevented from spreading |
| Test | 2019 Now I have to tell you a nice story. |
| 2020 Now I have to tell you a nice story. | |
| 2021 Now I have to tell you a nice story. | |
| 2022 Now I have to tell you a nice story. | |
| 2023 Now I have to tell you a nice story. |
| Tag | Train | Test |
| 2018 | unseen | 0.736 |
| 2019 | 0.721 | 0.737 |
| 2020 | 0.735 | 0.744 |
| 2021 | 0.749 | 0.749 |
| 2022 | unseen | 0.747 |
| 2023 | unseen | 0.747 |
| 2024 | unseen | 0.739 |
| 2025 | unseen | 0.747 |
One would expect the metric to produce the same score based solely on the translation quality. However, as listed in Table 9, we prove that by merely changing the year tag, we can influence the average score of the test set. During training, the model would be able to observe that 2019 is associated with the worst score and 2021 the best. During test time, the model follows this bias and also extrapolates it to upcoming years where it predicts an improvement in the average DA scores. While the differences might appear small, they are on the same scale as the differences between years in the training data.
Implication.
Our year-as-a-domain setting might be overly simple, but the vulnerability of COMET to latent biases cannot be neglected. Although Amrhein and Sennrich (2022) has shown that COMET is not sensitive to numbers, this work reveals that it can be systematically exploited in an artificial setting. We offer a more practical (adversarial) example that one may disguise biomedical domain translations as news translations to game COMET.
3.7 Lack of multi-reference support [usage]
In machine translation, there usually exist many valid translations for the same input. An effective metric should incorporate multiple ground truths if available, thereby enhancing the accuracy and robustness of its evaluation. Existing metrics like BLEU or chrF rely on surface-level overlap to capture the ground truth space from multiple references while metrics like ParBLEU (Bawden et al., 2020) can automatically generate paraphrases of a given reference to be included during evaluation.
By design, only one reference can be used as input to the COMET architecture. Whilst one may argue that representing a text reference in a neural space can ease the restriction on word choices, it might still be beneficial to use multiple references to overcome defects in the base embedding model. Therefore, we are curious whether COMET can explicitly and reliably leverage multiple references. We identify three distinct ways in which multiple references have been incorporated in COMET in previous literature (Rei et al., 2020b; Zouhar and Bojar, 2024):
max: Taking the maximum over the scores from multiple passes with different references.
avg: Averaging the scores from multiple passes with different references.
agg: Obtaining an aggregate score per example as follows. A quadruplet of source , hypothesis , reference , and alternative reference is fed to COMET six times in different [src, hyp, ref] arrangements: , as well as . Then, the average score from these six passes is multiplied by where denotes the standard deviation. See Rei et al. (2020b) for further explanation.
| WMT23 | COMET | |||||
| ref | ref (alt) | avg | max | agg | ||
| MQM | HeEn | 0.885 | 0.897 | 0.910 | 0.910 | 0.949 |
| EnDe | 0.974 | 0.936 | 0.974 | 0.987 | 0.974 | |
| ZhEn | 0.783 | 0.908 | 0.850 | 0.858 | 0.950 | |
| DA | EnDe | 0.885 | 0.949 | 0.910 | 0.897 | 0.910 |
| ZhEn | 0.717 | 0.875 | 0.783 | 0.775 | 0.883 | |
| DeEn | 0.901 | 0.912 | 0.923 | 0.923 | 0.912 | |
| EnZh | 0.933 | 0.817 | 0.900 | 0.867 | 0.933 | |
| CsUk | 0.846 | 0.802 | 0.901 | 0.890 | 0.868 | |
| EnCs | 0.858 | 0.875 | 0.858 | 0.858 | 0.867 | |
| EnJa | 0.941 | 0.824 | 0.934 | 0.934 | 0.926 | |
| JaEn | 0.922 | 0.915 | 0.928 | 0.928 | 0.922 | |
We use additional references from the WMT23 test set if available (HeEn) or the outputs from the best-scoring system in each direction in the metrics shared task (Freitag et al., 2023) as an alternative reference. We report pairwise system-level accuracy (Kocmi et al., 2021) for various translation directions in Table 10. Our results suggest that there is no single method that can consistently take advantage of the inclusion of multiple references with the existing COMET implementation. At a higher inference cost, the six-pass aggregation with COMET might have a tiny edge over other methods when MQM is treated as human ground truths, but it is also outperformed under DA by single-reference or other multi-reference methods.
As translation systems have greatly improved lately, the above pattern might be explained by Freitag et al. (2020)’s finding that high-quality translation outputs do not benefit from multi-reference evaluation. We also caution that these observations are highly dependent on the quality of the underlying references. As studied previously, obtaining high-quality references is not trivial (Freitag et al., 2020, 2023; Zouhar and Bojar, 2024). Our use of the top-performing system outputs as alternate references is fit for the purpose but not optimal.
Recommendation.
Our recommendations for the inclusion of multiple references into COMET or even other neural metrics are aspirational as this topic warrants further investigation. Extending unified pre-training (Wan et al., 2022) with multiple references in the architecture as well as using training objectives more suitable for handling more than one references (Zheng et al., 2018; Fomicheva et al., 2020) can be helpful.
3.8 Translationese [data]
COMET has been trained with human translations as references and machine translations as hypotheses, where both could be considered “translationese” to a certain extent (Gellerstam, 1986). An intriguing question is whether COMET would be affected by translationese and whether it prefers a translationese to a more natural translation.
| WMT19 EnDe | COMET | Ranking | COMET | Ranking | ||||
| orig. hyp | para. hyp | orig. hyp | para. hyp | orig. hyp | para. hyp | orig. hyp | para. hyp | |
| Sys-6823 | 0.845 | 0.840 | 8 | 5 | 0.8207 | 0.8243 | 8 | 4 |
| Sys-6862 | 0.867 | 0.842 | 2 | 2 | 0.8397 | 0.8281 | 2 | 3 |
| Sys-6819 | 0.843 | 0.835 | 9 | 9 | 0.815 | 0.8171 | 9 | 8 |
| Sys-6808 | 0.869 | 0.843 | 1 | 1 | 0.8402 | 0.8284 | 1 | 2 |
| Sys-6974 | 0.866 | 0.842 | 3 | 3 | 0.8384 | 0.8286 | 3 | 1 |
| Sys-6651 | 0.847 | 0.835 | 7 | 8 | 0.8214 | 0.8166 | 7 | 9 |
| Sys-6926 | 0.852 | 0.839 | 4 | 6 | 0.8243 | 0.8218 | 4 | 6 |
| Sys-6763 | 0.851 | 0.841 | 5 | 4 | 0.8229 | 0.8237 | 5 | 5 |
| Online-B | 0.838 | 0.830 | 10 | 10 | 0.8046 | 0.8095 | 10 | 10 |
| Sys-6731 | 0.850 | 0.837 | 6 | 7 | 0.8215 | 0.8204 | 6 | 7 |
Translationese in references.
We first conduct an experiment to see if the translationese present in the reference would undermine system evaluation with COMET. We consider WMT’s official reference as a standard version and Freitag et al. (2020)’s paraphrased reference as a less translationese reference.444The “paraphrased as-much-as-possible” version. Experiments are carried out under two settings: 1) WMT19 EnDe submissions evaluated by COMET and 2) WMT20 EnDe evaluated by COMET (Rei et al., 2020b). These two settings cover two scenarios—whether the test suite has been used in training the COMET model, or not. A breakdown of COMET scores and rankings for individual systems are listed in Appendix C Tables 13 and 14.
We notice that COMET scores decline dramatically when we switch the reference from the original one to the paraphrased one—aiming to reduce translationese. It means that COMET is indeed sensitive to such changes. Yet interestingly, the overall system ranking in either setting remains rather stable. We find a very high Kendall’s of 0.9827 and 0.9833 on the system rankings in the two settings; pairwise accuracy computed against human judgements also maintained at 0.924. We conclude that translationese in the reference impacts absolute COMET scores but not system ranking. These patterns are consistent regardless of whether the model has been exposed to the test set.
[style=mdframedlisting]
Translationese in hypotheses.
We then attempt to understand if a varying degree of translationese in the system outputs will influence system ranking by COMET. We run COMET and COMET on WMT19 EnDe system outputs as well as their corresponding rephrased outputs against the same source and reference. To acquire paraphrases affordably, we shortlist the top-10 systems’ translations in the previous experiment and we prompt GPT-4o using the prompt outlined in Figure 3.555Accessed gpt-4o-2024-08-06 via API in Aug 2024. We do not feed the source sentence to prevent the model from revising the quality.
We observe that COMET and COMET yield the same system ranking when the original hypotheses are scored. After substituting the hypotheses with their paraphrases, the ranking has changed more under COMET which witnesses much lower pairwise accuracy and Kendall’s compared to COMET. This suggests that COMET is more sensitive to potential changes in the degree of translationese than COMET, but we are unable to measure accuracy or correlation w.r.t. human judgements due to the lack of such.
Considerations.
We note the limitations in the experimental design: 1) we assume that the paraphrased references/hypotheses are as good as the original ones and less translationese, which might be hard to verify; 2) evaluations that anchor to human judgements assume that human evaluators provide assessment solely on the quality (and do not overly insist on adequacy/translationese); 3) our comparison between COMET and COMET only shows that they do not behave the same in dealing with change in translationese without informing us which is more satisfactory. Investigating the above problems would be costly but interesting.
3.9 Model reporting [usage]
Different COMET models can yield distinct results. Therefore it is important to always specify the specific model for sensible score interpretation and comparison. In this section, we look at to what extent this holds up in scientific literature.
We automate this bibliometric task with SemanticScholar API (Kinney et al., 2023).
Starting with 1.1K papers that cite one of the COMET papers, we obtain 417 papers from 2021 to 2024 that have an easily accessible PDF version.666Papers in 2020 did not have to report the specific model version as there was only one available at the time.
We parse the PDF text and check if any of the tables contains the string comet.
Within those papers, we check whether the COMET model information is contained in the PDF using a regular expression.777At least one of the following: xcomet\-;
wmt(20|21|22|23)\-comet; wmt\-da\-estimator;
comet[ \-](da|20|21|22|23).
After further manual validation, we found that 50 of the examined papers do not report a specific COMET version.
This establishes a lower bound that at least 12% of papers report COMET scores without specific model information.
In addition, out of the almost 1000 papers running COMET in their evaluation, most only cite the first COMET paper (Rei et al., 2020a) instead of the paper that describes the specific models that are being used (Rei et al., 2020b, 2022b, 2022a, 2022c, 2023a, 2023b; Glushkova et al., 2021; Wan et al., 2022; Alves et al., 2024; Guerreiro et al., 2023).
Recommendation.
Always report the specific COMET model, ideally with a link. In addition, it is better to cite the affiliated COMET paper as opposed to the first paper (Rei et al., 2020a), because different checkpoints have variations in training regimes that might be crucial in analysing the evaluation outcome.
3.10 Discussions on other issues
Interpretation of significance.
A statistical hypothesis test merely shows how likely the difference between the average of a model’s scores on a test set and that of another model’s scores on the same test set is caused by random fluctuation. As Kocmi et al. (2024b) examined, the significance between two metric scores can be made arbitrarily high and one can force by using a sizeable test set. However, this tells little about whether this difference is meaningful to a human reader. For this reason, we stress the use of mt-thresholds that converts differences in metric scores to how perceivable they are by human annotators.888kocmitom.github.io/MT-Thresholds
Averaging and subtracting COMET scores.
Research nowadays favours experiments on multiple translation directions, as multilingual translation models and large language models become trendy. Recent papers are more often seen to report a (macro-)average COMET score as an aggregate measure to reflect model performance, often in XxEn, EnXx, and AllAll. Whilst indicative to some extent, this custom is not entirely scientific because 1) the score range is inherently distinct for each translation direction and 2) there is no assurance that the scores are on a linear scale. Consequently, an outlying score in a single direction can distort the average, leading to a false claim. Likewise, absolute COMET score differences are not comparable if from different base numbers or directions. We suggest that, when multiple translation directions are of interest, it might be more viable to report the number of wins (against another system) as an aggregation of individual scores.
Optimizing towards COMET.
Owing to its strong correlation with human judgement, there have been recent works that investigate the feasibility of using COMET to further improve translation quality. These strategies either include COMET in a distillation workflow Finkelstein and Freitag (2024); Guttmann et al. (2024), as a data filtering method Peter et al. (2023), as a decoding method Freitag et al. (2022a); Fernandes et al. (2022); Vernikos and Popescu-Belis (2024) or as a training objective Yan et al. (2023).
Nevertheless, as a neural metric, COMET may cease to be a good measure if practitioners over-optimize a system towards it. As Yan et al. (2023) demonstrated, a model trained towards COMET can generate “universal translations” (hallucinations) preferred by COMET regardless of the source sentence. Recently, using COMET-based MBR decoding has become prevalent in shared tasks, e.g. at the IWSLT24 offline translation task by the CMU team. Their system has the highest COMET scores but ranks third in terms of human evaluation in EnDe Ahmad et al. (2024). Similarly, Unbabel-Tower70B at WMT24 also used MBR and dominated all automatic metrics at WMT24 (Kocmi et al., 2024a). We await WMT24’s human evaluation results to understand the implications further.
MBR decoding could be seen as an automatic way to exploit bias in the scoring method, (currently COMET in most cases), so practitioners need to beware of its shortcomings, disclose the use of such, and base system building on multiple (less correlated) metrics (Jon et al., 2023).
Sensitivity to sentence segmentation.
Like most MT metrics, COMET is a sentence-oriented metric, but sometimes in the evaluation of MT, we do not provide the systems with sentence-segmented input. This is most often the case in speech translation (ST) where sentence segmentation is treated as part of the task Ahmad et al. (2024). To address the problem of mismatching segmentation between the system output and the reference, a common solution in ST is to re-segment the output using a minimum error rate method (Matusov et al., 2005) in order to force-align it with the reference. Forced alignment can introduce segmentation errors resulting in truncated (thus grammatically incorrect) sentences. There is evidence that COMET, as a metric reliant on sentence embeddings, is more sensitive to these segmentation errors than metrics like BLEU, which rely purely on -grams with no linguistic notion of the sentence (Amrhein and Haddow, 2022). However in their comparison of COMET with human ranking, Sperber et al. (2024) suggested that COMET-based ranking is robust to segmentation errors yet that a “more thorough study of this issue is needed”.
4 The SacreCOMET Package
To help alleviate problems 3.1 (software version), 3.2 (compute precision), and 3.9 (model reporting), we release a simple package sacrecomet with two functionalities. Given a model name, the first functionality attempts to find the appropriate citation including a link to the paper and a BibTeX: {mdframed} [style=mdframedlisting] ⬇ $ pip install sacrecomet $ sacrecomet cite Unbabel/xcomet-xl https://arxiv.org/abs/2310.10482 @misc{guerreiro2023xcomet, title={xCOMET: Transparent Machine Translation Evaluation through Fine-grained Error Detection}, ... The second functionality semi-automatically detects the local software versions to generate a signature for better reproducibility. Both functionalities can also be run in interactive mode. {mdframed}[style=mdframedlisting]
5 Conclusion
COMET is currently one of the most powerful automatic metrics or quality estimators for machine translation, consistently achieving the top correlation with human judgement. In comparison to previous metrics, it is a statistical learning model and thus inherits all the related problems in addition to possible technical misconfigurations. We urge practitioners to consider more deeply the use of COMET in non-standard scenarios especially where such training bias might come into play. Beyond these issues, there has been confusion in the literature in reporting the correct COMET model and its correct setting. For improved consistency, we release an easy-to-use tool to assist practitioners.
Acknowledgments
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant numbers 10052546 and 10039436].
Appendix A Batch size and GPU/CPU
We run the test inference on a combination of GPU or CPU with varying batch sizes (BS, 1 or 100). Results in Table 12 demonstrate that the tiny effects of these choices are negligible for COMET reporting.
| Difference | MAE | ||
| BS=1, GPU | BS=1, GPU | ||
| BS=1, GPU | BS=64 GPU | ||
| BS=1, GPU | BS=1, CPU | ||
| BS=64, GPU | BS=64, CPU | ||
Appendix B Distribution of COMET scores for empty and valid hypothesis
Appendix C System rankings with paraphrased reference
| WMT20 EnDe | COMET | Ranking | ||
| original ref | paraphrased ref | original ref | paraphrased ref | |
| Sys-1069 | 0.508 | 0.305 | 9 | 9 |
| Sys-832 | 0.540 | 0.333 | 8 | 8 |
| Sys-1535 | 0.560 | 0.356 | 5 | 5 |
| Online-A | 0.499 | 0.288 | 10 | 10 |
| Online-B | 0.554 | 0.351 | 7 | 6 |
| Online-G | 0.268 | 0.052 | 14 | 14 |
| Online-Z | 0.329 | 0.121 | 13 | 13 |
| Sys-73 | 0.405 | 0.192 | 12 | 12 |
| Sys-1520 | 0.563 | 0.360 | 4 | 4 |
| Sys-890 | 0.578 | 0.371 | 3 | 3 |
| Sys-1136 | 0.472 | 0.264 | 11 | 11 |
| Sys-388 | 0.102 | -0.074 | 16 | 16 |
| Sys-737 | 0.555 | 0.351 | 6 | 7 |
| Ref-A | 0.878 | 0.526 | 1 | 1 |
| Ref-B | 0.591 | 0.446 | 2 | 2 |
| Sys-179 | 0.189 | 0.000 | 15 | 15 |
| WMT19 EnDe | COMET | Ranking | ||
| original ref | paraphrased ref | original ref | paraphrased ref | |
| Sys-6862 | 0.867 | 0.817 | 2 | 2 |
| Sys-6820 | 0.834 | 0.780 | 12 | 12 |
| Sys-6819 | 0.843 | 0.789 | 9 | 9 |
| Sys-6651 | 0.847 | 0.794 | 7 | 7 |
| Sys-6926 | 0.852 | 0.797 | 4 | 5 |
| Sys-6808 | 0.869 | 0.818 | 1 | 1 |
| Sys-6785 | 0.837 | 0.785 | 11 | 11 |
| Sys-6974 | 0.866 | 0.814 | 3 | 3 |
| Sys-6763 | 0.851 | 0.797 | 5 | 6 |
| Sys-6674 | 0.811 | 0.756 | 17 | 17 |
| Sys-6508 | 0.804 | 0.752 | 19 | 19 |
| Sys-6731 | 0.850 | 0.797 | 6 | 4 |
| Sys-6871 | 0.809 | 0.756 | 18 | 18 |
| Sys-6479 | 0.833 | 0.777 | 13 | 13 |
| Sys-6823 | 0.845 | 0.792 | 8 | 8 |
| Sys-6790 | 0.386 | 0.364 | 22 | 22 |
| Sys-6981 | 0.826 | 0.774 | 14 | 14 |
| Online-A | 0.815 | 0.761 | 16 | 16 |
| Online-B | 0.838 | 0.784 | 10 | 10 |
| Online-G | 0.795 | 0.738 | 20 | 20 |
| Online-X | 0.728 | 0.673 | 21 | 21 |
| Online-Y | 0.821 | 0.762 | 15 | 15 |